Find the Transitive closure by using Warshall’s Algorithm where A= {1, 2, 3,
4, 5, 6} and R= {(x,y)| (x-y)=2}
1
Expert's answer
2020-11-25T15:57:30-0500

Here the steps of Warshall’s Algorithm:


Step 1: Execute "W:=M_R, k:=0"


Step 2: Execute "k:=k+1"


Step 3: For all "i\ne k" such that "w_{ik}=1", and for all "j" execute the operation "w_{ij}:=w_{ij}\lor w_{kj}"


Step 4: If "k=n", then stop: the solution is "W=M_{R^*}", else go to step 2.


In our case "A= \{1, 2, 3,4, 5, 6\}" and "R= \{(x,y)\ |\ x-y=2\}".


So "n=6" and "R=\{(3,1),(4,2),(5,3),(6,4)\}".



"W^{(0)}=M_R=\left[\begin{array}{cccccc} 
0 & 0 & 0 & 0 & 0 & 0\\
0 & 0 & 0 & 0 & 0 & 0\\
1 & 0 & 0 & 0 & 0 & 0\\
0 & 1 & 0 & 0 & 0 & 0\\
0 & 0 & 1 & 0 & 0 & 0\\
0 & 0 & 0 & 1 & 0 & 0
\end{array}\right]"


"W^{(1)}=\left[\begin{array}{cccccc} 
0 & 0 & 0 & 0 & 0 & 0\\
0 & 0 & 0 & 0 & 0 & 0\\
1 & 0 & 0 & 0 & 0 & 0\\
0 & 1 & 0 & 0 & 0 & 0\\
0 & 0 & 1 & 0 & 0 & 0\\
0 & 0 & 0 & 1 & 0 & 0
\end{array}\right]"


"W^{(2)}=\left[\begin{array}{cccccc} 
0 & 0 & 0 & 0 & 0 & 0\\
0 & 0 & 0 & 0 & 0 & 0\\
1 & 0 & 0 & 0 & 0 & 0\\
0 & 1 & 0 & 0 & 0 & 0\\
0 & 0 & 1 & 0 & 0 & 0\\
0 & 0 & 0 & 1 & 0 & 0
\end{array}\right]"


"W^{(3)}=\left[\begin{array}{cccccc} 
0 & 0 & 0 & 0 & 0 & 0\\
0 & 0 & 0 & 0 & 0 & 0\\
1 & 0 & 0 & 0 & 0 & 0\\
0 & 1 & 0 & 0 & 0 & 0\\
1 & 0 & 1 & 0 & 0 & 0\\
0 & 0 & 0 & 1 & 0 & 0
\end{array}\right]"


"W^{(4)}=\left[\begin{array}{cccccc} 
0 & 0 & 0 & 0 & 0 & 0\\
0 & 0 & 0 & 0 & 0 & 0\\
1 & 0 & 0 & 0 & 0 & 0\\
0 & 1 & 0 & 0 & 0 & 0\\
1 & 0 & 1 & 0 & 0 & 0\\
0 & 1 & 0 & 1 & 0 & 0
\end{array}\right]"


"W^{(5)}=\left[\begin{array}{cccccc} 
0 & 0 & 0 & 0 & 0 & 0\\
0 & 0 & 0 & 0 & 0 & 0\\
1 & 0 & 0 & 0 & 0 & 0\\
0 & 1 & 0 & 0 & 0 & 0\\
1 & 0 & 1 & 0 & 0 & 0\\
0 & 1 & 0 & 1 & 0 & 0
\end{array}\right]"


"M_{R^*}=W^{(6)}=\left[\begin{array}{cccccc} 
0 & 0 & 0 & 0 & 0 & 0\\
0 & 0 & 0 & 0 & 0 & 0\\
1 & 0 & 0 & 0 & 0 & 0\\
0 & 1 & 0 & 0 & 0 & 0\\
1 & 0 & 1 & 0 & 0 & 0\\
0 & 1 & 0 & 1 & 0 & 0
\end{array}\right]"


"R^*=\{(3,1),(4,2),(5,1), (5,3),(6,2),(6,4)\}".